08-09 October 2008 Rotator repair.

Observers: D. Thompson, J.Hill (Tucson), LUCIFER folks (N.Ageorges,W.Seifert, et al)
Software Support: C.Biddick, D. Cox in Tucson
Telescope Operators: D.Gonzalez-Huerta, W.Wack

Summary

Rotator cable chain failed at beginning of the night. Lost the entire night aside from some flexure calibration measurements on LUCIFER.

Details

LUCIFER evening twilight flats? No, calibration unit and flexure compensation calibration work first.

Note: Startup problems: we had a power outage at some point during the day. Recovery included having to re-issue the "selectAGW" command, and a button in the treehouse needed to be pushed to reset something for the rotator. These should be documented as part of a normal operational startup checklist for the OSAs to work through. It would also be a good idea to document a proper startup in terms of collimation: clear M1 & M2 active optics, focus using Z global offset on M1, record offset and mirror temp, then correct pointing.

Diagnosing rotator cable chain problems occupied 7:30 - 11:30 MST, after which the conclusion was to bring up a spare part and an engineer from Tucson tomorrow.

We experimented briefly with manual guiding and IDL wavefront sensing without the rotator, but this was fairly tedious and only about 80% successful. The rest of the night was spent on off-sky work with LUCIFER calibration lamp. Of course, the seeing was good and the sky was clear.

-- DavidThompson - 08 Oct 2008
Topic revision: r4 - 09 Oct 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback